Ask
your own question, for FREE!
Mathematics
12 Online
If I pick 5 letter of the alphabet and I chose another person to pick 6 letter from the alphabet. If any of mine match the other persons letters I win. I none match I lose. Why is there a 72 percent chance of me winning?
Still Need Help?
Join the QuestionCove community and study together with friends!
compute the probability that the other person wins she has to pick letter, without replacement that do not match your five say you pick a, b, c, d, e then the probability she doesn't pick your letters is \[\frac{21}{26}\times \frac{20}{25}\times \frac{19}{24}\times \frac{18}{23}\times \frac{17}{22}\times \frac{16}{21}\]
^nice. P(you win) = 1 - P(other person wins)
could also say i guess that the probability the other person wins is \[\binom{21}{26}\] same thing i suppose
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
VERNEISH:
Two objects are moving along separate linear paths where each path is described by position, d, and time, t.
Maci:
what are the best stock apps for people who want to start investing?
Twaylor:
genuienly how do i get better self control i keep watcing youtube and theres this
IzzyCookie:
|dw:1792074056138:dw| Uhm, I don't know where to ask this, but I'm in college welding at my community college, and I need help learning vertical welding for
thehearken:
not sure if this is considered spam, im sorry if it is, but does anyone have tips
3 hours ago
1 Reply
0 Medals
4 days ago
3 Replies
2 Medals
1 week ago
3 Replies
0 Medals
2 days ago
9 Replies
1 Medal
1 week ago
2 Replies
1 Medal